# Interactive Program Synthesis for Modeling Collaborative Physical Activities from Narrated Demonstrations

### v1

@@ Abstract.
Teaching systems physical tasks is a long-standing goal in HCI, yet most prior work has focused on non-collaborative physical activities. Collaborative tasks introduce added complexity, requiring systems to infer users’ assumptions about their teammates’ intent—an inherently ambiguous and dynamic process. This necessitates representations that are interpretable and correctable, enabling users to inspect and refine system behavior. We address this challenge by framing collaborative task learning as a program synthesis problem. Our system represents behavior as editable programs and uses narrated demonstrations—paired physical actions and natural language—as a unified modality for teaching, inspecting, and correcting system logic without requiring users to see or write code. The same modality is used for the system to communicate its learning to users. In a within-subjects study, 20 users taught multiplayer soccer tactics to our system. 70% (14/20) of participants successfully refined learned programs to match their intent and 90% found it easy to correct the programs. The study surfaced unique challenges in representing learning as programs and in enabling users to teach collaborative physical activities; we discuss these issues and outline mitigation strategies.
program synthesis, interactive programming, probabilistic programming language, machine learning, artificial intelligence, human computer interactions
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; .; .isbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. Our system learns collaborative physical activities from narrated mixed reality demonstrations, synthesizes programs with domain-specific APIs, and demonstrates back to the user for feedback and refinement.:
@@ 1. Introduction
However, realizing such a system presents significant technical challenges. First, learning collaborative activities entails substantially greater complexity compared to single-user tasks.. In HCI, systems for learning physical tasks have predominantly relied on demonstrations and, more recently, natural language. While teaching via demonstrations can capture the nuances of physical movement, it often fails to infer the underlying logic or user intent (Cypher, 1993). Conversely, natural language can articulate high-level goals but struggle to ground abstract commands into specific, continuous trajectories in the physical world—a classic instance of the symbol grounding problem (Harnad, 1990; Bisk et al., 2020). Collaboration further increases this complexity since user’s assumptions on collaborators’ intent also need to be accounted for. Also, the constraint that the system must be teachable by users without a technical background necessitates that its internal models of physical activities be both interpretable and directly editable through an interface.
@@
The existing literature on systems for representing collaborative physical activities is sparse. The most closely related works, such as GhostAR (Cao et al., 2019) and VisCourt (Cheng et al., 2024), leverage mixed reality (MR) to generate virtual avatars for demonstrating interactions. This approach provides the system with a fully observable view of the behaviors of, and interactions between, the user and their collaborators. However, rather than teaching collaborative tasks, users configure the interactions to generate tutorials. For example, in VisCourt (Cheng et al., 2024), a basketball coach selects and parameterizes pre-defined tactical plays from a menu-based interface. In our problem of interest, we aim to have users teach collaborative activities to the system using natural modalities (e.g., language and demonstrations).
@@
We present an interactive system that enables non-programmers to model collaborative physical activities by synthesizing programs from narrated demonstrations. Our choice of this input modality—physical demonstrations paired with concurrent natural language—is informed by prior work showing that it facilitates more accurate and robust learning (Tung et al., 2018; Waytowich et al., 2019; Yu and Mooney, 2023); narration disambiguates the intent behind an action, while the demonstration grounds language in the physical world. As summarized in Fig. 1, our system treats narrated demonstration as a key interaction primitive across the entire teaching loop—for initial instruction, inspection, and correction—enabling users to align the program’s behavior with their intent without interacting with raw source code. The system allows a user to first teach a task via a narrated demonstration in mixed reality (MR). The system processes this recording with a large vision-language model (LVLM) to generate an initial program. The system communicates its learning via decision flow diagram (visualizing the underlying code structure) and program execution which provides narrated demonstrations to the user. To correct the program, the user provides feedback by providing narration along with physical demonstration via annotations.
@@
    An interactive programming synthesis system for modeling collaborative physical activities that enables:
@@
      non-programmers to teach and correct the system via narrated demonstrations, and
@@
      the system to effectively communicate its learning to users via narrated demonstrations.
@@
    A user study demonstrating that non-programmers can use the proposed interactive framework to:
    1. (a)
      easily teach, inspect, and correct the system so as to model complex collaborative physical activities as programs, and
    2. (b)
      align program behaviors with the intent of users
@@ 2.1. Tutorial Authoring Systems
Augmented Reality (AR) has long been envisioned for delivering tutorials and task guidance. Seminal systems demonstrated knowledge-based overlays for step guidance (Feiner et al., 1993) and proactive, sensor-driven instructions for physical assembly (Antifakos et al., 2002). However, these early systems focused on the delivery of guidance, with authoring relegated to experts rather than end users. A significant thread in HCI research has focused on empowering practitioners to author their own tutorials. For example, InstruMentAR (Liu et al., 2023b) and AdapTutAR (Huang et al., 2021) enable experts to record embodied demonstrations to produce step-wise AR overlays for machine operation, with AdapTutAR further tailoring guidance to the learner’s state. Beyond AR/VR, the community has explored auto-authoring from demonstrations in mixed media. TutoriVR (Kumaravel et al., 2019) embeds 2D videos within VR workspaces accompanied by 3D contextual cues, while VAAR (Yamaguchi et al., 2020) reconstructs a 3D assembly graph from 2D video to align AR animations. TutorialLens (Kong et al., 2021) advances this by directly authoring AR tutorials from synchronized narration and demonstration, capturing finger trajectories and speech to structure the steps.
A critical distinction in recent work is the role of natural language. While much prior work simply records and replays narration to augment steps, newer systems analyze language to infer procedural structure. For instance, Truong et al. automatically derive two-level hierarchical procedures from instructional videos (Truong et al., 2021), and TutoAI (Chen et al., 2024) systematizes pipelines to extract tutorial components from videos and transcripts to assemble mixed-media guides. Complementing these authoring tools, recent work demonstrates fully automated AR guidance from Large Language and Vision Models (LLVMs), enriching steps with embedded visual cues (Zhao et al., 2025). While this highlights the increasing integration of modern AI, its focus is on automated generation rather than end-user teaching and refinement.
Despite these advances, most systems target single-user, non-collaborative tasks. They seldom model role-dependent logic, joint timing, or the state of teammates and opponents. While systems like GhostAR (Cao et al., 2019) and VisCourt (Cheng et al., 2024) explore multi-agent choreography and tactic authoring in AR and MR, they primarily focus on configuring pre-defined interactions rather than learning an editable, closed-loop policy from narrated demonstrations.
@@ 2.2. Interactive Task Learning
In Interactive Task Learning (ITL), the goal is for a machine to learn task knowledge from a human teacher, often incrementally and through interactions that build an interpretable representation (Laird et al., 2017). Prominent HCI examples include ONYX (Ruoff et al., 2023), which enables users to teach natural-language interfaces by aligning speech and demonstrations with system capabilities, and VAL (Lawley and MacLellan, 2024), which uses LLMs in a neuro-symbolic pipeline to acquire editable, hierarchical procedures from conversation. These systems excel at software-centric or simulated tasks, emphasizing symbolically interpretable knowledge and conversational repair (e.g., clarifications, confirmations, undo). However, they typically do not ground learning in continuous physical trajectories or reason about multi-agent coordination and timing—challenges central to collaborative physical activities.
@@ 2.3. Interactive Program Synthesis
A parallel research thread focuses on synthesizing executable programs from end-user input, such as examples or demonstrations. Canonical work like FlashFill (Gulwani, 2011) demonstrated how carefully designed domain-specific languages (DSLs) could yield powerful transformations from a few examples. Subsequent HCI research has developed interaction models to make synthesized programs more understandable and editable. This includes techniques for disambiguating synthesized programs via navigation and active-example queries (Mayer et al., 2015), and augmented-example workflows that allow users to iteratively repair programs (Zhang et al., 2020).
Domain-specific systems show how end users can create and refine automations involving complex logic like loops and conditionals. Examples include Rousillon (Chasins et al., 2018) for hierarchical web scraping, DiLogics (Pu et al., 2023) for handling diverse input conditions, and MIWA (Chen et al., 2023) for mixed-initiative diagnosis and repair. In mobile computing, SUGILITE (Li et al., 2017) and PUMICE (Li et al., 2019) combine natural language with programming-by-demonstration (PbD) (Cypher, 1993) to synthesize procedures that users can inspect and modify. In data science, Wrex (Drosos et al., 2020) produces readable notebook code from examples, Falx (Wang et al., 2021) synthesizes visualization pipelines from example outputs, and recent interfaces like SQLucid (Tian et al., 2024) generate step-by-step, editable natural-language explanations to help non-experts understand and correct model-generated SQL.
@@ 3. Background
3.1. Large Vision Language Models (LVLMs)
Large Vision Language Models extend large language models by processing visual inputs in addition to linguistic ones. LVLMs have demonstrated strong capabilities in aligning natural language descriptions with structured visual data, supporting tasks such as captioning, multimodal retrieval, and program synthesis (Radford et al., 2021; Yuan et al., 2021; Li et al., 2022). Their capacity to integrate symbolic and perceptual representations makes them well suited for interpreting narrated demonstrations, which combine visual demonstrations of physical activity with verbal explanations. In our work, LVLMs serve as the program synthesis tool.
@@ 3.2. Scenic Probabilistic Programming Language
Scenic is a probabilistic programming language designed for modeling complex physical environments with stochastic variation (Fremont et al., 2019, 2023)). By allowing users to specify distributions over 3-dimensional scene configurations and object behaviors, Scenic enables flexible yet interpretable representations of physical contexts. This stochasticity is particularly valuable for collaborative physical activities, where environmental conditions and agent interactions naturally vary across instances.
@@ 4. System Architecture
4.1. Teaching Interface in Mixed Reality
Environment Modeling. To model and generate MR environments involving multiple entities with interactive behaviors, we use Scenic (Fremont et al., 2019), a probabilistic programming language designed to specify distributions of physical scenarios. The probabilistic aspect of the language will later be utilized to model variations in user’s narrated demonstrations. In this system, a Scenic program is provided, which models the initial scenes and the behaviors of environment agents, i.e. the entities that the user will interact with. [^1] It would be ideal if users can flexibly author the environments they desire from scratch. This authoring may be possible provided that user can teach just one’s own behavior, which we aim to examine through our user study. The full authoring is out of scope of this paper.
Action Recognition. Having large vision language models to understand sophisticated human actions, especially collaborative physical activities, is still an open challenge. Thus, we restrict the user actions to be triggered via MR controller buttons such that the system can easily recognize and log the user’s actions.
Embodied Narrated Demonstrations. User wears a MR headset which generates a pre-scripted environment overlaid on the physical world. The user is immersed in the environment physically interacting with environment agents. From this embodied perspective, the user provides physical demonstrations, interacting with environment agents, while verbally explaining the rationale behind actions. During the narrated demonstrations, the MR headset records the following data: (1) RGB video of the demonstrations recorded from virtual camera, (2) timestamped annotation of actions (e.g., [12:33] “User passed to Teammate”), (3) transcription of narration which consists of timestamped tokens (e.g., [12:34] “Therefore”, [12:35] “you”).

### v2

# Interactive Program Generation for Modeling Collaborative Physical Activities from Narrated Demonstrations

@@ Abstract.
Teaching systems physical tasks is a long-standing goal in HCI, yet most prior work has focused on non-collaborative physical activities. Collaborative tasks introduce added complexity, requiring systems to infer users’ assumptions about their teammates’ intent—an inherently ambiguous and dynamic process. This necessitates representations that are interpretable and correctable, enabling users to inspect and refine system behavior. We address this challenge by framing collaborative task learning as a program generation problem. Our system represents behavior as editable programs and uses narrated demonstrations—paired physical actions and natural language—as a unified modality for teaching, inspecting, and correcting system logic without requiring users to see or write code. The same modality is used for the system’s responses, enabling intuitive two-way interaction. In a within-subjects study, 20 users taught multiplayer soccer tactics to our system. Participants successfully refined learned programs to match their intent and found the system easy to use.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. Our system learns collaborative physical activities from narrated mixed reality demonstrations, generates programs with domain-specific APIs, and demonstrates back to the user for feedback and refinement.:
@@ 1. Introduction
However, realizing such a system presents significant technical challenges. First, learning collaborative activities entails substantially greater complexity compared to single-user tasks.. In HCI, systems for learning physical tasks have predominantly relied on demonstrations and, more recently, natural language. While teaching via demonstrations can capture the nuances of physical movement, it often fails to infer the underlying logic or user intent (cypher1993watch). Conversely, natural language can articulate high-level goals but struggle to ground abstract commands into specific, continuous trajectories in the physical world—a classic instance of the symbol grounding problem (harnad1990symbol; bisk2020experience). Collaboration further increases this complexity since user’s assumptions on collaborators’ intent also need to be accounted for. Also, the constraint that the system must be teachable by users without a technical background necessitates that its internal models of physical activities be both interpretable and directly editable through an interface.
@@
The existing literature on systems for representing collaborative physical activities is sparse. The most closely related works, such as GhostAR (cao2019ghostar) and VisCourt (Cheng2024VisCourt), leverage mixed reality (MR) to generate virtual avatars for demonstrating interactions. This approach provides the system with a fully observable view of the behaviors of, and interactions between, the user and their collaborators. However, rather than teaching collaborative tasks, users configure the interactions to generate tutorials. For example, in VisCourt (Cheng2024VisCourt), a basketball coach selects and parameterizes pre-defined tactical plays from a menu-based interface. In our problem of interest, we aim to have users teach collaborative activities to the system using natural modalities (e.g., language and demonstrations).
@@
We present an interactive system that enables non-programmers to model collaborative physical activities by synthesizing programs from narrated demonstrations. Our choice of this input modality—physical demonstrations paired with concurrent natural language—is informed by prior work showing that it facilitates more accurate and robust learning (tung2018narrated; waytowich2019narration; yu2023deltaco); narration disambiguates the intent behind an action, while the demonstration grounds language in the physical world. As summarized in Fig. 1, our system treats narrated demonstration as a key interaction primitive across the entire teaching loop—for initial instruction, inspection, and correction—enabling users to align the program’s behavior with their intent without interacting with raw source code. The system allows a user to first teach a task via a narrated demonstration in mixed reality (MR). The system processes this recording with a large vision-language model (LVLM) to generate an initial program. The system communicates its learning via decision flow diagram (visualizing the underlying code structure) and program execution which provides narrated demonstrations to the user. To correct the program, the user provides feedback by providing narration along with physical demonstration via annotations.
@@
    An interactive programming system for modeling collaborative physical activities that uses narrated demonstrations as a unified modality for:
@@
      non-programmers to teach and correct the system, and
@@
      the system to effectively communicate its learning to users.
@@
    A user study in a tactical soccer setting, together with an existence proof in a factory environment, demonstrating the feasibility of our system for modeling collaborative physical activities as users intend.
@@ 2.1. Tutorial Authoring Systems
Augmented Reality (AR) has long been envisioned for delivering tutorials and task guidance. Seminal systems demonstrated knowledge-based overlays for step guidance (Feiner1993KnowledgeAR) and proactive, sensor-driven instructions for physical assembly (Antifakos2002Proactive). However, these early systems focused on the delivery of guidance, with authoring relegated to experts rather than end users. A significant thread in HCI research has focused on empowering practitioners to author their own tutorials. For example, InstruMentAR (Liu2023InstruMentAR) and AdapTutAR (Huang2021AdapTutAR) enable experts to record embodied demonstrations to produce step-wise AR overlays for machine operation, with AdapTutAR further tailoring guidance to the learner’s state. Beyond AR/VR, the community has explored auto-authoring from demonstrations in mixed media. TutoriVR (Kumaravel2019TutoriVR) embeds 2D videos within VR workspaces accompanied by 3D contextual cues, while VAAR (Yamaguchi2020VAAR) reconstructs a 3D assembly graph from 2D video to align AR animations. TutorialLens (Kong2021TutorialLens) advances this by directly authoring AR tutorials from synchronized narration and demonstration, capturing finger trajectories and speech to structure the steps.
A critical distinction in recent work is the role of natural language. While much prior work simply records and replays narration to augment steps, newer systems analyze language to infer procedural structure. For instance, Truong et al. automatically derive two-level hierarchical procedures from instructional videos (Truong2021Auto), and TutoAI (Chen2024TutoAI) systematizes pipelines to extract tutorial components from videos and transcripts to assemble mixed-media guides. Complementing these authoring tools, recent work demonstrates fully automated AR guidance from Large Language and Vision Models (LLVMs), enriching steps with embedded visual cues (Zhao2025GuidedReality). While this highlights the increasing integration of modern AI, its focus is on automated generation rather than end-user teaching and refinement.
Despite these advances, most systems target single-user, non-collaborative tasks. They seldom model role-dependent logic, joint timing, or the state of teammates and opponents. While systems like GhostAR (cao2019ghostar) and VisCourt (Cheng2024VisCourt) explore multi-agent choreography and tactic authoring in AR and MR, they primarily focus on configuring pre-defined interactions rather than learning an editable, closed-loop policy from narrated demonstrations.
@@ 2.2. Interactive Task Learning
In Interactive Task Learning (ITL), the goal is for a machine to learn task knowledge from a human teacher, often incrementally and through interactions that build an interpretable representation (Laird2017ITL). Prominent HCI examples include ONYX (Ruoff2023ONYX), which enables users to teach natural-language interfaces by aligning speech and demonstrations with system capabilities, and VAL (Lawley2024VAL), which uses LLMs in a neuro-symbolic pipeline to acquire editable, hierarchical procedures from conversation. These systems excel at software-centric or simulated tasks, emphasizing symbolically interpretable knowledge and conversational repair (e.g., clarifications, confirmations, undo). However, they typically do not ground learning in continuous physical trajectories or reason about multi-agent coordination and timing—challenges central to collaborative physical activities.
2.3. Task Learning with Feedback in Reinforcement Learning.
A substantial body of reinforcement learning (RL) research studies how agents can learn tasks from explicit human feedback rather than hand-crafted reward functions. Surveys of interactive RL and human-in-the-loop methods catalog a wide range of feedback modalities, including scalar evaluative rewards, critiques, demonstrations, and natural-language guidance (Najar2021HumanAdviceSurvey). Early work such as the TAMER framework treats human input as real-valued evaluative feedback delivered while the agent acts; the agent learns a reward model directly from this feedback and optimizes its policy accordingly (KnoxStone2008TAMER). Deep TAMER extends this idea to high-dimensional visual domains using deep networks, showing that sparse human reward signals can successfully shape Atari agents (Warnell2018DeepTAMER). Preference-based RL instead elicits pairwise comparisons over trajectory segments, training a reward model that explains human preferences and then optimizing the policy with respect to that learned reward (Christiano2017HumanPreferences). More recent interactive imitation-learning methods combine multiple forms of feedback—corrective demonstrations, scalar evaluations, and implicit positive feedback—and use uncertainty estimates to actively query humans when clarification is most useful (Celemin2023CorrectiveEvaluative). Together, this literature demonstrates that RL agents can leverage diverse feedback channels (evaluative rewards, preferences, and corrections) to learn complex behaviors with relatively few interactions, but typically focuses on single-agent tasks with relatively sparse literature in the space of collaborative physical activities.
2.4. Programming by Demonstration with Corrective Feedback.
Robot Programming by Demonstration (PbD), or Learning from Demonstration (LfD), aims to let users specify robot behaviors by showing examples rather than writing code (Billard2008PbDHandbook; Argall2009LfDSurvey). Surveys and tutorials describe a spectrum of approaches that encode demonstrations as trajectories, movement primitives, finite-state machines, or probabilistic models, and highlight design choices in how demonstrations are collected (kinesthetic teaching, teleoperation, motion capture) and generalized to new situations (ChernovaThomaz2014RLFromTeachers; Calinon2019LfDEncyclopedia). While many PbD systems treat demonstrations as fixed training data, some explicitly incorporate user feedback to refine learned programs. For instance, Argall et al. allow users to provide corrective feedback on low-level motion primitives, using this input to scaffold and refine policies built from previously demonstrated skills (Argall2011TeacherFeedback). Such work shows that combining demonstration with post-hoc corrections can improve performance and data efficiency, but typically offers only limited support for inspecting or editing the learned task representation itself (e.g., at the level of high-level decision logic). Our work builds on these ideas by representing tasks as interpretable programs and designing interfaces that treat feedback, inspection, and correction as interaction mechanisms for modeling collaborative physical activities.
@@ 2.5. Interactive Program Synthesis
A parallel research thread focuses on synthesizing executable programs from end-user input, such as examples or demonstrations. Canonical work like FlashFill (Gulwani2011FlashFill) demonstrated how carefully designed domain-specific languages (DSLs) could yield powerful transformations from a few examples. Subsequent HCI research has developed interaction models to make synthesized programs more understandable and editable. This includes techniques for disambiguating synthesized programs via navigation and active-example queries (Mayer2015DisambPBE), and augmented-example workflows that allow users to iteratively repair programs (Zhang2020AugEx). Prior work also shows that explicitly surfacing symbolic representations of program structure helps non-programmers inspect logic and make targeted edits, for example by exposing block-level units of generated code (yen2024coladder; jiang2022genline), visualizing and monitoring execution traces of LLM-generated code (xie2024waitgpt), or decomposing analysis tasks into editable subgoals that users can steer and verify (kazemitabaar2024interactive).
Domain-specific systems show how end users can create and refine automations involving complex logic like loops and conditionals. Examples include Rousillon (Chasins2018Rousillon) for hierarchical web scraping, DiLogics (Pu2023DiLogics) for handling diverse input conditions, and MIWA (Chen2023MIWA) for mixed-initiative diagnosis and repair. In mobile computing, SUGILITE (Li2017SUGILITE) and PUMICE (Li2019PUMICE) combine natural language with programming-by-demonstration (PbD) (cypher1993watch) to synthesize procedures that users can inspect and modify. In data science, Wrex (Drosos2020Wrex) produces readable notebook code from examples, Falx (Wang2021Falx) synthesizes visualization pipelines from example outputs, and recent interfaces like SQLucid (Tian2024SQLucid) generate step-by-step, editable natural-language explanations to help non-experts understand and correct model-generated SQL.
@@ 3. Background
3.1. Large Vision Language Models
Large Vision Language Models (LVLMs) extend large language models by processing visual inputs in addition to linguistic ones. LVLMs have demonstrated strong capabilities in aligning natural language descriptions with structured visual data, supporting tasks such as captioning, multimodal retrieval, and program synthesis (radford2021clip; yuan2021florence; li2022glip). Their capacity to integrate symbolic and perceptual representations makes them well suited for interpreting narrated demonstrations, which combine visual demonstrations of physical activity with verbal explanations. In our work, LVLMs serve as the program synthesis tool.
@@ 3.2. Scenic Probabilistic Programming Language
Scenic (fremont2019scenic; scenic-mlj23) is a domain-specific probabilistic programming language for modeling and generating physical environments with stochastic multi-agent behaviors and initial conditions in simulation. A snippet of a Scenic program is shown in Fig. 4. When compiling such a program, Scenic imports (i) the map of the simulated world, (ii) definitions of objects (e.g., ball, goal) and agents (e.g., Coach, Teammate), and (iii) an API library defining the action space (e.g., MoveTo, Pass, Speak) and support functions (e.g., HasBallPossession()). Using this information, the environment is modeled as follows. Lines 22–28 specify a distribution over initial conditions by instantiating the user and other agents with initial positions and assigned behaviors. Because the world map has been imported, coordinates are interpreted relative to the map origin. The user’s behavior (CoachBehavior()) is modeled using the provided APIs. At execution time, Scenic samples an initial condition for each agent: the user’s location is fixed at the map origin, while the teammate’s and opponent’s positions are sampled from specified ranges relative to the user’s position and orientation. After instantiating objects and agents, Scenic executes each agent’s behavior function to generate a dynamic, reactive environment.
@@ 4. System Architecture
4.1. Teaching Interface in Mixed Reality
Environment Modeling. To model and generate MR environments involving multiple entities with interactive behaviors, we use Scenic (fremont2019scenic), a probabilistic programming language designed to specify distributions of physical scenarios.
Action Space. We assume that the action space is provided. These actions (e.g., dribbling, kicking) can be executed by either users physically moving or using MR controllers. For example, the user may physically move around to re-position oneself, but for kicking a virtual ball, one needs to use the controller to execute this action.
Embodied Narrated Demonstrations. The user wears a mixed reality (MR) headset and is immersed in a 3D workspace, embodying the specific agent they aim to teach. While embodying this agent, the user authors the initial scene and assigns behaviors to surrounding agents to incrementally build a dynamic environment. In the running soccer example, the user first constructs the scene by placing a teammate, an opponent, goalposts, and a ball at desired locations via verbal instructions and MR controller input (e.g., “Place my teammate here” [user points at (12.3, 4.26)]). The user then verbally assigns behaviors to these surrounding agents (e.g., “The teammate should pass the ball to me.”). A large language model is used in real time to map verbal instructions and controller annotations to actions from a provided action space, identify which agent or object each action targets, and immediately execute the corresponding actions to generate the instructed environment. As the user incrementally builds this dynamic environment, one provides narrated demonstrations of how the focal agent should react to and coordinate with surrounding agents. Unlike surrounding agents, the user can enact the focal agent’s behavior through physical demonstrations while simultaneously narrating and annotating via the controllers. During this teaching phase, the MR headset records: (1) RGB video of the user and surrounding agents and objects from a virtual camera, (2) timestamped annotations of actions triggered with MR controllers (e.g., [12:33] User passed to Teammate”), and (3) a transcription of the narration as timestamped tokens (e.g., [12:34] Therefore”, [12:35] “you”).